OpenStudy (88):

Solve. 4 5/6+w+2 3/4=12 A.5 5/12 B.5 1/5 C.4 7/12 D.4 5/12

OpenStudy (88):

help plz

OpenStudy (88):

wrong

OpenStudy (88):

can someone help me plz like @ganeshie8 @LegendaryClasher @Mhudson00

OpenStudy (88):

@cmtboy2

OpenStudy (88):

help plz

pooja195 (pooja195):

Is 4 part of the question ?

OpenStudy (88):

no

pooja195 (pooja195):

So it looks like this? \[\huge~\rm~\frac{ 5 }{ 6}+w+2\frac{ 3 }{ 4 }=12\]

OpenStudy (88):

no

pooja195 (pooja195):

Can you draw it out?

OpenStudy (88):